OpenStudy (anonymous):

The current in a simple electrical circuit is inversely proportional to the resistance. If the current is 20 amperes ( an ampere is a unit for measuring current) when the resistance is 3 ohms, find the current when the resistance is 7.4 ohms.

OpenStudy (anonymous):

8.1081 amperes

OpenStudy (anonymous):

how do you write that as a proper fraction?

OpenStudy (anonymous):

\[I _{1}R _{1}=I _{2}R _{2}\]\[20*3=I _{2}*7.4\]\[I _{2}=8.1081\]

OpenStudy (anonymous):

It says to write is as proper fraction or mixed number?

OpenStudy (anonymous):

you can write the answer \[I _{2}=8.1\]

OpenStudy (anonymous):

I tried it said it was wrong and i have to type it in fraction form :/

OpenStudy (anonymous):

it shows it like this...

OpenStudy (anonymous):

\[y\left(\begin{matrix}x \\ k\end{matrix}\right)\]

OpenStudy (anonymous):

what is that?

OpenStudy (anonymous):

I don't understand

OpenStudy (anonymous):

an improper fraction sorry it looks different idk how to show it.

OpenStudy (anonymous):

That was just example... but 8.1081 is supposed to be turned into a improper fraction

OpenStudy (anonymous):

300/37

OpenStudy (anonymous):

alright i will try that.

OpenStudy (anonymous):

I got it right thanks

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

just put 3/8